A sleep researcher hypothesizes that sleep is related to feelings of tiredness. Below are partial SPSS results with α = 0.01 of the analysis for a sample of participants at a sleep center.

fs_corr_exm_app_pkg_pc_fall_01_b_img

a) Compute the appropriate test statistic(s) to make a decision about H0.
(Hint: Make sure to write down the null and alternative hypotheses to help solve the problem.)
Critical value =  ; Test statistic =
Decision:  ---Select--- Reject H0 Fail to reject H0

b) Compute the corresponding effect size(s) and indicate magnitude(s).
If not appropriate, input and/or select "na" below.
Effect size =  ;   ---Select--- na trivial effect small effect medium effect large effect

c) Make an interpretation based on the results.

More sleep is related to feeling more tired.More sleep is related to feeling less tired.    Sleep is not related to feeling tired.
Answer #1

The null and alternative hypothesis is, Ho: p=0 H: P70 Degrees of freedom=n-2= 13-2=11 Critical value oft at 0.01 level of si